OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (lncognlto):

m = y2-y1/x2-x1 = (5-6)/(5+1) = -1/6 therefore m of the perpendicular = 6 as m x m(perp) = -1 mid point = (x1+x2)/2, (y1+y2)/2 = 4/2, 11/2 = 2, 11/2 subst in y = mx + c 11/2 = 6(2) + c c= -13/2 therefore y = 6x - 13/2 or 2y = 12x - 13

OpenStudy (anonymous):

what is it in standard form

OpenStudy (lncognlto):

According to www.mathsisfun.com/algebra/standard-form.html, the standard form for a linear equation is Ax + By = C. So here it would be 12x - 2y = 13.

OpenStudy (anonymous):

goood looking thanks

OpenStudy (lncognlto):

You're welcome! ;)